OpenStudy (anonymous):

I just need to make sure that I am correct on this. Would 180 miles per 2 hours be converted to 264 feet per second?

OpenStudy (phi):

change miles to feet: \[ 180 \cancel{\text{ miles}} \cdot \frac{5280 \text{ feet}}{1 \cancel{\text{ miles}}} \] change hours to seconds \[ 2 \cancel{\text{ hours}} \cdot \frac{ 60 \cancel{\text{ min}}}{1 \cancel{\text{ hours}} } \cdot \frac{ 60\text{ sec}}{1 \cancel{\text{ min}} }\]

OpenStudy (phi):

then divide the feet by the seconds to get ft/sec

OpenStudy (anonymous):

I am not quite understanding

OpenStudy (phi):

do you know how to convert miles to feet? every mile is 5280 feet that means 1 mile is 5280 feet 2 miles is twice as far or 2*5280 feet 3 miles is three times 5280 feet and so on what is 180 miles in feet ?

OpenStudy (anonymous):

those are the conversions from miles to feet, and from hours to seconds

OpenStudy (anonymous):

from miles per hour to feet per second

OpenStudy (anonymous):

180 miles in feet is 950,400 feet

OpenStudy (phi):

next, can you change 2 hours to seconds ?

OpenStudy (anonymous):

yes. 7200 seconds. right?

OpenStudy (phi):

now you know \[ \frac{180 \text{ miles}}{2 \text{ hours}} = \frac{ 950400 \text{ ft}}{7200 \text{ sec}}\]

OpenStudy (anonymous):

so, now do I just divide 950,400 by 7200

OpenStudy (phi):

yes, you should simplify that fraction

OpenStudy (anonymous):

so it would simplify to 132 ft per sec.

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (phi):

we could have made the problem a little bit easier by immediately simplifying 180 miles/ 2 hours to 90 miles / 1 hour and then changing 90 miles to feet, and 1 hour to 3600 seconds

OpenStudy (anonymous):

yeah. but my teacher wants us to set it up like this 

OpenStudy (phi):

yes, but that does not prevent you from simplifying 180/2 into 90/1

OpenStudy (anonymous):

I understand.

OpenStudy (phi):

and if I were doing the conversion, I might write it like you just did. But it is more difficult to explain in these tiny boxes, so I did the top and bottom separately.

OpenStudy (anonymous):

ohhh okay.

OpenStudy (anonymous):

And I have another question that says "tickets cost $19.95 each, and there is a $3 charge for each order no matter how many tickets are ordered. Write an expression for the cost in dollars of ordering tickets." would I put 19.95x + 3 ?

OpenStudy (anonymous):

or 19.95x + 3x ?

OpenStudy (phi):

It might help your thinking to define what "x" represents. x is the number of tickets. the cost for ordering "x" tickets will be x * cost of each ticket plus $3 charge in other words x*19.95 + 3 which is usually written as 19.95x + 3 (we put the number in front of the variable)